OpenStudy (anonymous):

True or False, cos(x/5)sin(x/5)=1/2 sin(2x/5) I think it is True.

OpenStudy (aum):

sin(2A) = 2sin(A)cos(A) sin(A)cos(A) = 1/2 * sin(2A). Therefore, cos(x/5)sin(x/5)=1/2 sin(2x/5)

OpenStudy (anonymous):

So it is true?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

great can you help with another one?